(3) (4 pts) Nitrous acid will not disproportionate spontaneously to the gases nitrogen dioxide and nitric oxide. Show this by writing the balanced chemical reaction and evaluating the standard emf (in V) for the reaction.
Answer #1

The balanced chemical equation for disproportionation reaction is shown below.

2HNO3(aq) = NO2(aq) + NO + H2O(1)

Two half cell reactions are shown below.

H NO2(aq) + H+ (aq) + + NO(g) + H2O(1) Ered = +1.00 V HNO3(aq) + NO3(aq) + H+ (aq) + Eorid = -1.12 V

Ecell = Ered + Eorid Ecell = 1.00 V +(-1.12 V) El= -0.12 V

Negative value of E^o_{cell} indicates non spontaneous reaction.